# Resonances and resonance expansions for $\delta$-interactions on the half-space

### v2

@@ 5.1. Dirichlet Boundary Condition
$\frac{\left\lvert\ln{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}+i\frac{\pi}{2}\right\rvert}{\ln{\Bigl{(}h^{-\beta}\Bigr{)}}\left\lvert 8\pi y_{3}+(2k+1)i\pi h^{\beta}\right\rvert}\leq\frac{1}{h^{\beta}\ln{(h^{-\beta})}}\Biggl{(}\left\lvert\frac{\ln{\left\lvert\ln{(w)}+(2k+1)i\pi\right\rvert}}{\ln{(w)}+(2k+1)\pi}\right\rvert+\frac{\frac{\pi}{2}}{\left\lvert{\ln{(w)}+(2k+1)\pi}\right\rvert}\Biggr{)}.$
@@
(5.18) $\left\lvert\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{h^{\beta}\ln{\Bigl{(}h^{-\beta}\Bigr{)}}}{4\pi y_{3}}.$
By Eq. 5.3, we can also write
(5.19) $\left\lvert R_{k}-\frac{\ln{\Bigl{(}\ln{(w)}+(2k+1)i\pi\Bigr{)}}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{h^{2\beta}\ln^{2}{\Bigl{(}h^{-\beta}\Bigr{)}}}{8\pi^{2}y_{3}^{2}}.$
Turning back to Eq. 5.17, we now have
@@
This expression can be also bounded using $\left\lvert\operatorname{Arg}{\biggl{(}8\pi y_{3}h^{-\beta}+(2k+1)i\pi\biggr{)}}\right\rvert\leq\frac{\pi}{2}$, $\ln{(1+t)}\leq t$, the expression for $t$, $(2k+1)^{2}\leq 2(4k^{2}+1)$ and the second inequality in Eq. 5.9 as
@@
For $\beta>1$, the dominant term for small $h$ are the first and last term. More specifically the first is dominant for $\beta\geq 2$ and the other for $21<\beta<$ 2. So, for $h$ small enough
@@
Now we look at the real part of Eq. 5.16
@@
Using that $\left\lvert\operatorname{Arg}{(1+is)}\right\rvert\leq 2\left\lvert s\right\rvert$ for small $s$, the second inequality in Eq. 5.9, Eq. 5.18 and Eq. 5.19 we get the bound
(5.20) $\left\lvert\operatorname{Re}{z_{k}}+\frac{(2k+1)\pi h}{2y_{3}}\right\rvert\leq\frac{\varepsilon^{-1}h^{\beta}}{2\pi y_{3}}+\frac{h^{\beta+1}}{8y_{3}^{2}}+\frac{h^{2\beta+1}\ln^{2}{\Bigl{(}h^{-\beta}\Bigr{)}}}{16\pi^{2}y_{3}^{3}}+\frac{h^{\beta+1}\ln{\Bigl{(}h^{-\beta}\Bigr{)}}}{8\pi y_{3}^{2}}\leq\frac{\varepsilon^{-1}h^{\beta}}{\pi y_{3}},$
@@
By factorizing, Eq. 5.20 and by the second inequalities in Eq. 5.9 and Eq. 5.8 respectively
@@
Now we observe that $(2k+1)^{2}\geq k^{2}$ $\forall k\in\mathbb{Z}$, $k\neq 0$ (which is escluded by Eq. 5.8). Using this fact, together with the reciprocal of the first inequality of Eq. 5.9, we have
@@
To get Eq. 5.14 and Eq. 5.15 we use respectively that $\forall k\in\mathbb{Z}$, $k\neq 0$, $(2k+1)^{2}\leq 16k^{2}$ and that $(2k+1)^{2}\geq k^{2}$, together with Eq. 5.9. ∎
Now we consider $\beta>1$ for Eq. 5.1-.
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq. 5.1- satisfying
(5.21) $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$
@@
which, using Eq. 5.9, implies
@@
Since $\beta>1$, by the second inequality in Eq. 5.9, we have
@@
for $h$ small enough. Recalling Eq. 5.7, we write
@@
(5.22) $\begin{cases}\operatorname{Re}{z_{k}}=\frac{h}{2y_{3}}\Biggr{(}-2k\pi+\operatorname{Arg}{\biggr{(}1-\frac{i(2k+1)}{8y_{3}}h^{\beta}\biggr{)}}-\operatorname{Im}{R_{k}}\Biggr{)}\\ \operatorname{Im}{z_{k}}=\frac{h}{2y_{3}}\Biggr{(}-\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}-\ln{\left\lvert 1-\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}+\operatorname{Re}{R_{k}}\Biggr{)}\end{cases}.$
@@
Eq. 5.18 and Eq. 5.19 are true and can be proved as done in Proposition 5.2. So, by Eq. 5.22
@@
This expression can be also bounded using $\left\lvert\operatorname{Arg}{\biggl{(}8\pi y_{3}h^{-\beta}+(2k+1)i\pi\biggr{)}}\right\rvert\leq\pi$, $\ln{(1+t)}\leq t$, the expression for $t$, $(2k+1)^{2}\leq 2(4k^{2}+1)$ and the second inequality in Eq. 5.9 as
@@
For $\beta>1$, the dominant term for small $h$ are the first and last term. More specifically the first is dominant for $\beta\geq 2$ and the other for $21<\beta<$ 2. What said implies that for $h$ small enough
@@
Now we look at the real part of Eq. 5.22
@@
Using that $\left\lvert\operatorname{Arg}{(1+is)}\right\rvert\leq 2\left\lvert s\right\rvert$ for small $s$, the second inequality in Eq. 5.9, Eq. 5.18 and Eq. 5.19 we get the bound
(5.23) $\left\lvert\operatorname{Re}{z_{k}}+\frac{k\pi h}{y_{3}}\right\rvert\leq\frac{\varepsilon^{-1}h^{\beta}}{2\pi y_{3}}+\frac{h^{\beta+1}}{8y_{3}^{2}}+\frac{h^{2\beta+1}\ln^{2}{\Bigl{(}h^{-\beta}\Bigr{)}}}{16\pi^{2}y_{3}^{3}}+\frac{h^{\beta+1}\ln{\Bigl{(}h^{-\beta}\Bigr{)}}}{8\pi y_{3}^{2}}\leq\frac{\varepsilon^{-1}h^{\beta}}{\pi y_{3}},$
@@
This new inequality can be used to bound the quantity $B=\left\lvert\frac{y_{3}}{2k^{2}\pi^{2}h}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\biggl{(}\operatorname{Re}{z_{k}}\biggr{)}^{2}-\frac{h}{2y_{3}}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\right\rvert$. By factorizing, Eq. 5.23 and by the inequalities Eq. 5.9 and Eq. 5.21
@@
$\left\lvert\operatorname{Im}{z_{k}+\frac{y_{3}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}}{2k^{2}\pi^{2}h}\biggl{(}\operatorname{Re}{z_{k}}\biggr{)}^{2}}\right\rvert$ $\leq\left\lvert\operatorname{Im}{z_{k}}+\frac{h}{2y_{3}}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\right\rvert+\left\lvert\frac{2y_{3}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}}{h(2k+1)^{2}\pi^{2}}\biggl{(}\operatorname{Re}{z_{k}}\biggr{)}^{2}-\frac{h}{2y_{3}}\ln{\Bigl{(}8\pi y_{3}h^{-\beta}\Bigr{)}}\right\rvert$
@@ 5.2. Neumann Boundary Condition
By the definition of $-h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{N}}$ it follows that the resonances are the solutions of
(5.24) ±$\pm h^{-\beta}-\frac{iz}{4\pi h}-\frac{e^{2iy_{3}\frac{z}{h}}}{8\pi y_{3}}=0,$
which can be rewritten as $w_{\pm}=x_{\pm}e^{x_{\pm}}$, where $x_{\pm}=\pm 8\pi y_{3}h^{-\beta}-2iy_{3}\frac{z}{h}$ and $w_{\pm}=e^{\pm 8\pi y_{3}h^{-\beta}}$. So the solutions are still expressed in terms of the Lambert W function, with the only difference of it having a positive argument this time. A corresponding asymptotic expansion is still valid (one has only to substitute $12k+$ 1 with $2k$, reason being the change in sign of $w$)
@@
(5.25) $R_{k}=\sum_{j=0}^{+\infty}\sum_{m=1}^{+\infty}c_{j,m}\frac{\ln^{m}{\Bigl{(}\ln{(w)}+2ik\pi\Bigr{)}}}{\Bigl{(}\ln{(w)}+2ik\pi\Bigr{)}^{j+m}}.$
@@
(5.26) $z_{k}^{\pm}=\frac{ih}{2y_{3}}\biggl{(}2ik\pi-\ln{\Bigl{(}\ln{(y_{\pm})}+2ik\pi\Bigr{)}}+R_{k}\biggr{)}.$
@@
The series Eq. 5.25 is absolutely convergent for $w$ large (or small) enough and $k\in\mathbb{Z}$. More precisely we have the tail estimate
@@
Let $\varepsilon\in(0,1)$ ve given. Then, for $k$ such that $z_{k}$ is given by Eq. 5.26 and $\varepsilon\leq\left\lvert z_{k}\right\rvert\leq 1/\varepsilon$ we have
(5.27) $\frac{\varepsilon}{2}\leq\frac{\left\lvert k\right\rvert\pi h}{y_{3}}\leq\frac{2}{\varepsilon}.$
@@
Let $\beta\in(0,1)$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions of Eq. 5.24 satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq. 5.24+ satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
which, using Eq. 5.27, implies
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq. 5.24- satisfying
@@
which, using Eq. 5.27, implies
@@ 6. Resonance Expansion for the Wave and Schrödinger Equation with Point Interaction
As regards the wave equation, a similar analysis in the presence of a smooth $L_{\textup{comp}}^{\infty}(\mathbb{R}^{n})$ potential on the whole space is given in [22] (pages 39-45 and 110-111). We recall that $L^{\infty}_{\textup{comp}}(\mathbb{R}^{n})$ is the space of measurable, essentially bounded functions with compact support. In the following Theorem we will treat the abstract wave equation with generator given by a point interaction in +$\mathbb{R}^{3}_{+}$ and initial data in $L^{2}_{\textup{comp}}(\mathbb{R}^{3}_{+})$ and $H^{1}_{\textup{comp}}(\mathbb{R}^{3}_{+})$, the space of square integrable function in +$\mathbb{R}^{3}_{+}$ with compact support and the corresponding Sobolev space of order one. The restriction to data with compact support is needed, as in the standard case, to tame the exponentially diverging behavior of resonance at spatial infinity.
@@
(6.1) $\begin{cases}\biggl{(}\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\alpha,y}^{\textup{D}}\biggr{)}w(t)=0\\ w(0)=w_{0}\in H_{\textup{comp}}^{1}(\mathbb{R}_{+}^{3})\\ \frac{\partial w}{\partial t}(0)=w_{1}\in L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})\end{cases}.$
Let $\Set{z_{j}^{2}}$ and $\lambda$ be respectively the set of resonances and the possible eigenvalue of $\Delta_{\alpha,y}^{\textup{D}}$, and $R(z):=\Bigl{(}-\Delta_{\alpha,y}^{\textup{D}}-z^{2}\Bigr{)}^{-1}$ Then, for any $A>0$,
@@
    (6.2) $f_{j}=-\operatorname{Res}_{z=z_{j}}{\biggl{(}iR(z)w_{1}+zR(z)w_{0}\biggr{)}};$
@@
    where the sum is finite, $f_{j}$ are still defined as in Eq. 6.2 and
@@
(6.3) $\left\lVert\rho(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\leq C(1+\left\lvert z\right\rvert)^{j-1}e^{T(\operatorname{Im}{z})_{-}},\ \ \ j=0,1,2$
@@
where $z^{2}$ is not a resonance and it is denoted $(x)_{-}=\max\Set{-x,0}$.
@@
(6.4) $\Set{z\in\mathbb{C}}{\operatorname{Im}{z}\geq-A-\delta\ln{(1+\left\lvert z\right\rvert)}}\ .$
As shown previously, all resonances are on the curve Eq. 3.1. This means that the curve
(6.5) $\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl{(}\frac{1}{2y_{3}\left\lvert\operatorname{Re}{z}\right\rvert}\Biggr{)}}\ \ \ \ \ \text{equivalent to}\ \ \ \ \left\lvert\operatorname{Re}{z}\right\rvert=\frac{1}{2y_{3}}e^{-2y_{3}\operatorname{Im}{z}}$
is above all resonances. The equation of the curve which delimits the region Eq. 6.4 can be written in explicit form as follows
(6.6) $\left\lvert\operatorname{Re}{z}\right\rvert=\sqrt{\biggl{(}e^{-\frac{\operatorname{Im}{z}+A}{\delta}}-1\biggr{)}^{2}-\Bigl{(}\operatorname{Im}{z}\Bigr{)}^{2}}.$
Both curves are symmetric with respect to the imaginary axis, so we limit ourself to $\operatorname{Re}{z}>0$. If it happens that for $\operatorname{Im}{z}$ large enough, curve Eq. 6.6 is above curve Eq. 6.5, then for $\operatorname{Im}{z}$ large enough all $z$ which generate a resonance and obey this constraint are below Eq. 6.6 and hence only a finite quantity of them is in the region described by Eq. 6.4. This scenario happens for $\frac{1}{\delta}>2y_{3}$ regardless of the value of $A$.
In order to prove Eq. 6.3 we use Eq. 2.1 and write
(6.7) $\left\lVert\rho(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\leq\left\lVert\rho(-\Delta^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\\ +\left\lvert\Gamma_{\alpha,y}^{\textup{D}}(z)\right\rvert^{-1}\left\lVert\rho G_{z,y}^{\textup{D}}(-\Delta^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}.$
@@
where $L>\operatorname{diam}{(\operatorname{supp}{(\rho)})}$ and that is also valid if we replace $-\Delta$ with $-\Delta^{\textup{D}}$. This directly bounds the first term in Eq. 6.7. It does the same for the second one if we observe that $G_{z,y}^{\textup{D}}$ is in $L_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$ and that outside of resonances $\left\lvert\Gamma_{\alpha,y}^{\textup{D}}(z)\right\rvert^{-1}$ is bounded. Hence Eq. 6.3 follows. ∎
Proof of Theorem 6.1.

### v3

@@ 5.1. Dirichlet Boundary Condition
[…]
@@
$\left\lvert\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{h^{\beta}\ln{\Bigl(h^{-\beta}\Bigr)}}{4\pi y_{3}}.$ (5.18)
By Eq.˜5.3, we can also write
$\left\lvert R_{k}-\frac{\ln{\Bigl(\ln{(w)}+(2k+1)i\pi\Bigr)}}{\ln{(w)}+(2k+1)i\pi}\right\rvert\leq\frac{h^{2\beta}\ln^{2}{\Bigl(h^{-\beta}\Bigr)}}{8\pi^{2}y_{3}^{2}}.$ (5.19)
Turning back to Eq.˜5.17, we now have
@@
This expression can be also bounded using $\left\lvert\operatorname{Arg}{\biggl(8\pi y_{3}h^{-\beta}+(2k+1)i\pi\biggr)}\right\rvert\leq\frac{\pi}{2}$, $\ln{(1+t)}\leq t$, the expression for $t$, $(2k+1)^{2}\leq 2(4k^{2}+1)$ and the second inequality in Eq.˜5.9 as
@@
For $\beta>1$, the dominant terms for small $h$ are the first and the last term. More specifically the first is dominant for $\beta\geq 2$ and the other for $1<\beta<2$. So, for $h$ small enough
@@
Now we look at the real part of Eq.˜5.16
@@
Using that $\left\lvert\operatorname{Arg}{(1+is)}\right\rvert\leq 2\left\lvert s\right\rvert$ for small $s$, the second inequality in Eq.˜5.9, Eq.˜5.18 and Eq.˜5.19 we get the bound
$\left\lvert\operatorname{Re}{z_{k}}+\frac{(2k+1)\pi h}{2y_{3}}\right\rvert\leq\frac{\varepsilon^{-1}h^{\beta}}{2\pi y_{3}}+\frac{h^{\beta+1}}{8y_{3}^{2}}+\frac{h^{2\beta+1}\ln^{2}{\Bigl(h^{-\beta}\Bigr)}}{16\pi^{2}y_{3}^{3}}+\frac{h^{\beta+1}\ln{\Bigl(h^{-\beta}\Bigr)}}{8\pi y_{3}^{2}}\leq\frac{\varepsilon^{-1}h^{\beta}}{\pi y_{3}},$ (5.20)
@@
By factorizing, Eq.˜5.20 and by the second inequalities in Eq.˜5.9 and Eq.˜5.8 respectively
@@
Now we observe that $(2k+1)^{2}\geq k^{2}$ for all $k\in\mathbb{Z}$. Using this fact, together with the reciprocal of the first inequality of Eq.˜5.9, we have
@@
To get Eq.˜5.14 and Eq.˜5.15 we use respectively that for all $k\in\mathbb{Z}$, $k\neq 0$, $k^{2}\leq(2k+1)^{2}\leq 16k^{2}$, together with Eq.˜5.9. ∎
Now we consider $\beta>1$ for Eq.˜5.1-.
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq.˜5.1- satisfying
$\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ (5.21)
@@
which, using Eq.˜5.9, implies
@@
Since $\beta>1$, by the second inequality in Eq.˜5.9, we have
@@
for $h$ small enough. Recalling Eq.˜5.7, we write
@@
$\begin{cases}\operatorname{Re}{z_{k}}=\frac{h}{2y_{3}}\Biggr(-2k\pi+\operatorname{Arg}{\biggr(1-\frac{i(2k+1)}{8y_{3}}h^{\beta}\biggr)}-\operatorname{Im}{R_{k}}\Biggr)\\ \operatorname{Im}{z_{k}}=\frac{h}{2y_{3}}\Biggr(-\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}-\ln{\left\lvert 1-\frac{i(2k+1)}{8y_{3}}h^{\beta}\right\rvert}+\operatorname{Re}{R_{k}}\Biggr)\end{cases}.$ (5.22)
@@
Eq.˜5.18 and Eq.˜5.19 are true and can be proved as done in Proposition˜5.2. So, by Eq.˜5.22
@@
This expression can be also bounded using $\left\lvert\operatorname{Arg}{\biggl(8\pi y_{3}h^{-\beta}+(2k+1)i\pi\biggr)}\right\rvert\leq\pi$, $\ln{(1+t)}\leq t$, the expression for $t$, $(2k+1)^{2}\leq 2(4k^{2}+1)$ and the second inequality in Eq.˜5.9 as
@@
For $\beta>1$, the dominant term for small $h$ are the first and last term. More specifically the first is dominant for $\beta\geq 2$ and the other for $1<\beta<2$. What said implies that for $h$ small enough
@@
Now we look at the real part of Eq.˜5.22
@@
Using that $\left\lvert\operatorname{Arg}{(1+is)}\right\rvert\leq 2\left\lvert s\right\rvert$ for small $s$, the second inequality in Eq.˜5.9, Eq.˜5.18 and Eq.˜5.19 we get the bound
$\left\lvert\operatorname{Re}{z_{k}}+\frac{k\pi h}{y_{3}}\right\rvert\leq\frac{\varepsilon^{-1}h^{\beta}}{2\pi y_{3}}+\frac{h^{\beta+1}}{8y_{3}^{2}}+\frac{h^{2\beta+1}\ln^{2}{\Bigl(h^{-\beta}\Bigr)}}{16\pi^{2}y_{3}^{3}}+\frac{h^{\beta+1}\ln{\Bigl(h^{-\beta}\Bigr)}}{8\pi y_{3}^{2}}\leq\frac{\varepsilon^{-1}h^{\beta}}{\pi y_{3}},$ (5.23)
@@
This new inequality can be used to bound the quantity
$B=\left\lvert\frac{y_{3}}{2k^{2}\pi^{2}h}\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}\biggl(\operatorname{Re}{z_{k}}\biggr)^{2}-\frac{h}{2y_{3}}\ln{\Bigl(8\pi y_{3}h^{-\beta}\Bigr)}\right\rvert$.
By factorizing, Eq.˜5.23 and by the inequalities Eq.˜5.9 and Eq.˜5.21
@@
[…]
@@ 5.2. Neumann Boundary Condition
By the definition of $h^{2}\Delta_{\pm h^{-\beta},y}^{\textup{N}}$ it follows that the resonances are the solutions of
$\pm h^{-\beta}-\frac{iz}{4\pi h}-\frac{e^{2iy_{3}\frac{z}{h}}}{8\pi y_{3}}=0,$ (5.24)
which can be rewritten as $w_{\pm}=x_{\pm}e^{x_{\pm}}$, where $x_{\pm}=\pm 8\pi y_{3}h^{-\beta}-2iy_{3}\frac{z}{h}$ and $w_{\pm}=e^{\pm 8\pi y_{3}h^{-\beta}}$. So the solutions are still expressed in terms of the Lambert W function, with the only difference of it having a positive argument this time. A corresponding asymptotic expansion is still valid (one has only to substitute $2k+1$ with $2k$, reason being the change in sign of $w$)
@@
$R_{k}=\sum_{j=0}^{+\infty}\sum_{m=1}^{+\infty}c_{j,m}\frac{\ln^{m}{\Bigl(\ln{(w)}+2ik\pi\Bigr)}}{\Bigl(\ln{(w)}+2ik\pi\Bigr)^{j+m}}.$ (5.25)
@@
$z_{k}^{\pm}=\frac{ih}{2y_{3}}\biggl(2ik\pi-\ln{\Bigl(\ln{(y_{\pm})}+2ik\pi\Bigr)}+R_{k}\biggr).$ (5.26)
@@
The series Eq.˜5.25 is absolutely convergent for $w$ large (or small) enough and $k\in\mathbb{Z}$. More precisely we have the tail estimate
@@
Let $\varepsilon\in(0,1)$ ve given. Then, for $k$ such that $z_{k}$ is given by Eq.˜5.26 and $\varepsilon\leq\left\lvert z_{k}\right\rvert\leq 1/\varepsilon$ we have
$\frac{\varepsilon}{2}\leq\frac{\left\lvert k\right\rvert\pi h}{y_{3}}\leq\frac{2}{\varepsilon}.$ (5.27)
@@
Let $\beta\in(0,1)$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions of Eq.˜5.24 satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq.˜5.24+ satisfying $\varepsilon\leq\left\lvert z\right\rvert\leq{1}/{\varepsilon}$ obey
@@
which, using Eq.˜5.27, implies
@@
Let $\beta>1$ and $\varepsilon\in(0,1)$ be given. Then there is $h_{0}>0$ such that, when $h\in(0,h_{0}]$, all solutions to Eq.˜5.24- satisfying
@@
which, using Eq.˜5.27, implies
@@ 6. Resonance Expansion for the wave and Schrödinger Equation with Point Interaction
As regards the wave equation, a similar analysis in the presence of a smooth $L_{\textup{comp}}^{\infty}(\mathbb{R}^{n})$ potential on the whole space is given in [28] (pages 39-45 and 110-111). We recall that $L^{\infty}_{\textup{comp}}(\mathbb{R}^{n})$ is the space of measurable, essentially bounded functions with compact support. In the following Theorem we will treat the abstract wave equation with generator given by a point interaction in $\mathbb{R}^{3}_{+}$ and initial data in $L^{2}_{\textup{comp}}(\mathbb{R}^{3}_{+})$ and $H^{1}_{\textup{comp}}(\mathbb{R}^{3}_{+})$, the space of square integrable function in $\mathbb{R}^{3}_{+}$ with compact support and the corresponding Sobolev space of order one. The restriction to data with compact support is needed, as in the standard case, to tame the exponentially diverging behavior of resonance at spatial infinity.
@@
$\begin{cases}\biggl(\frac{\partial^{2}}{\partial t^{2}}-\Delta_{\alpha,y}^{\textup{D}}\biggr)w(t)=0\\ w(0)=w_{0}\in H_{\textup{comp}}^{1}(\mathbb{R}_{+}^{3})\\ \frac{\partial w}{\partial t}(0)=w_{1}\in L_{\textup{comp}}^{2}(\mathbb{R}_{+}^{3})\end{cases}.$ (6.1)
Let $\Set{z_{j}}$ be respectively the set of resonances including the possible eigenvalue of $\Delta_{\alpha,y}^{\textup{D}}$, and $R(z):=\Bigl(-\Delta_{\alpha,y}^{\textup{D}}-z^{2}\Bigr)^{-1}$ Then, for any $A>0$,
@@
    $f_{j}=-\operatorname{Res}_{z=z_{j}}{\biggl(iR(z)w_{1}+zR(z)w_{0}\biggr)};$ (6.2)
@@
    where the sum is finite, $f_{j}$ are still defined as in Eq.˜6.2 and
@@
$\left\lVert\rho(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\leq C(1+\left\lvert z\right\rvert)^{j-1}e^{T(\operatorname{Im}{z})_{-}},\ \ \ j=0,1,2$ (6.3)
@@
where $z$ is not a resonance and it is denoted $(x)_{-}=\max\Set{-x,0}$.
@@
[…]
As shown previously, all resonances are on the curve Eq.˜3.4. This means that the curve
$\operatorname{Im}{z}=\frac{1}{2y_{3}}\ln{\Biggl(\frac{1}{2y_{3}\left\lvert\operatorname{Re}{z}\right\rvert}\Biggr)}\ \ \ \ \ \text{equivalent to}\ \ \ \ \left\lvert\operatorname{Re}{z}\right\rvert=\frac{1}{2y_{3}}e^{-2y_{3}\operatorname{Im}{z}}$ (6.5)
is above all resonances. The equation of the curve which delimits the region Eq.˜6.4 can be written in explicit form as follows
$\left\lvert\operatorname{Re}{z}\right\rvert=\sqrt{\biggl(e^{-\frac{\operatorname{Im}{z}+A}{\delta}}-1\biggr)^{2}-\Bigl(\operatorname{Im}{z}\Bigr)^{2}}.$ (6.6)
Both curves are symmetric with respect to the imaginary axis, so we limit ourself to $\operatorname{Re}{z}>0$. If it happens that for $\operatorname{Im}{z}$ large enough, curve Eq.˜6.6 is above curve Eq.˜6.5, then for $\operatorname{Im}{z}$ large enough all $z$ which generate a resonance and obey this constraint are below Eq.˜6.6 and hence only a finite quantity of them is in the region described by Eq.˜6.4. This scenario happens for $\frac{1}{\delta}>2y_{3}$ regardless of the value of $A$.
In order to prove Eq.˜6.3 we use Proposition˜2.1 and write
$\left\lVert\rho(-\Delta_{\alpha,y}^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\leq\left\lVert\rho(-\Delta^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}\\ +\left\lvert\Gamma_{\alpha,y}^{\textup{D}}(z)\right\rvert^{-1}\left\lVert\rho G_{z,y}^{\textup{D}}(-\Delta^{\textup{D}}-z^{2})^{-1}\rho\right\rVert_{L^{2}(K)\to H^{j}(K)}.$ (6.7)
@@
where $L>\operatorname{diam}{(\operatorname{supp}{(\rho)})}$ and that is also valid if we replace $\Delta$ with $\Delta^{\textup{D}}$. This directly bounds the first term in Eq.˜6.7. It does the same for the second one if we observe that $G_{z,y}^{\textup{D}}$ is in $L_{\textup{loc}}^{2}(\mathbb{R}_{+}^{3})$ and that outside of resonances $\left\lvert\Gamma_{\alpha,y}^{\textup{D}}(z)\right\rvert^{-1}$ is bounded. Hence Eq.˜6.3 follows. ∎
Proof of Theorem˜6.1.